# Not Just for Archiving: Provable Benefits of Reusing the Archive in Evolutionary Multi-objective Optimization

### v1

@@ Introduction
Refer to caption Refer to caption Refer to caption Refer to caption
(a) KP 100D (b) NK 100D (c) TSP 100D (d) QAP 100D
Refer to caption Refer to caption Refer to caption Refer to caption
(e) KP 200D (f) NK 200D (g) TSP 200D (h) QAP 200D
Refer to caption Refer to caption Refer to caption Refer to caption
(i) KP 500D (j) NK 500D (k) TSP 500D (l) QAP 500D
Figure 2: Non-dominated solutions obtained by SMS-EMOA-A (archive only to store non-dominated solutions) and SMS-EMOA-AR (archive with reuse) on the multi-objective 0/1 Knapsack (KP), NK-Landscape ,(NK), TSP and QAP problems, with 100 (top panel), 200 (middle panel), and 500 (bottom panel) decision variables respectively. The Knapsack and NK-Landscape are maximization problems, and the TSP and QAP are minimization problems.
@@
In this section, we conduct experiments on two artificial problems OneJumpZeroJump and OneJumpZeroJumpSS analyzed in this paper, and four well-known practical optimization problems: the multi-objective 0/1 knapsack (KP) (Teghem 1994), travelling salesman problem (TSP) (Ribeiro et al. 2002), quadratic assignment problem (QAP) (Knowles and Corne 2003a) and NK-landscapes (NK) (Aguirre and Tanaka 2004), following the settings in (Li et al. 2024). Table 2 presents the results of SMS-EMOA solving OneJumpZeroJumpSS with problem size $n\in\{15,20,25,30\}$, $k=3$, and $a=2$, and Table 3 presents the results of SMS-EMOA soving OneJumpZeroJump with problem size $n\in\{15,20,25,30\}$ and $k=2$. We present the average number of generations over $1000$ runs for three variants: 1) SMS-EMOA with a large population size $\mu=n-2k+5$, 2) SMS-EMOA with an archive storing all non-dominated solutions ($\mu=5$), and 3) SMS-EMOA with archive reuse ($\mu=5$). For SMS-EMOA-A, since its expected running time on OneJumpZeroJump is at least $n^{\Omega(n)}$, we set the maximum number of iterations to $1,000,000$. For problem sizes $n=\{20,25,30\}$, a significant proportion of runs exceed this limit, thus these cases are marked with “-”. Results show that reusing the archive significantly reduces running time compared to using an archive only to store non-dominated solutions, and using a large population size.
@@
As for the four practical problems, each problem is instantiated at three sizes: 100, 200, and 500 variables. We conduct experiments using two SMS-EMOA variants: 1) Archive without reuse, and 2) Archive with reuse. Both variants follow common parameter settings: 1) the population size is set to 100; 2) the parent selection uses uniform random selection, as in the above theorems; 3) the crossover and mutation operators are problem-specific. For pseudo-Boolean problems (KP, NK), we use the uniform crossover (Eiben and Smith 2015) with a rate of 1.0 and the bit-wise mutation (Eiben and Smith 2015) with a rate of $1/n$, where $n$ is the number of variables. For TSP, we use the order-based crossover and the 2-opt mutation (Eiben and Smith 2015). For QAP, we use the cycle-based crossover and the 2-swap mutation (Eiben and Smith 2015). For both TSP and QAP, the crossover rate is 1.0 and the mutation rate is 0.05, following the practice in (Li et al. 2024). The budget for the algorithms is 10,000,000 fitness evaluations and the number of independent runs is set to 30. Table 4 presents the result of a widely used indicator, Hypervolume (HV) (Zitzler and Thiele 1999) of the two SMS-EMOA variants on the four problems. To compute the HV reference point, we sample 100,000 random solutions from the decision space and use the worst objective values among the non-dominated ones, following (Li et al. 2024). Table 4 shows that reusing the archive can lead to a better mean HV across all problem instances, where the differences in 9 out of 12 problem instances are significant.
@@
Figure 2 presents the non-dominated solutions obtained by the SMS-EMOA without archive reuse (blue dot) and SMS-EMOA-RA with archive reuse (black cross) on the multi-objective 0/1 Knapsack (KP), NK-Landscape (NK), TSP and QAP problems, with 100 (top panel), 200 (middle panel), and 500 (bottom panel) decision variables respectively. As can be seen, SMS-EMOA-RA (reusing the archive) consistently outperforms or matches SMS-EMOA (not reusing the archive) across all problem instances. The advantage of archive reuse becomes more pronounced as the problem size increases, particularly on the TSP.
$\dagger$” indicates that the result of not reusing the archive is significantly different from that of reusing it, based on the Wilcoxon rank-sum test at the 95% confidence level. Problem SMS-EMOA-A SMS-EMOA-AR KP-100 2.068e+06 (2.5e+04) $\dagger$ 2.098e+06 (2.9e+04) KP-200 6.062e+06 (8.4e+04) $\dagger$ 6.213e+06 (4.5e+04) KP-500 3.731e+07 (6.4e+05) $\dagger$ 3.897e+07 (3.4e+05) NK-100 5.986e-02 (3.9e-03) 6.194e-02 (2.1e-03) NK-200 5.584e-02 (2.7e-03) $\dagger$ 5.907e-02 (2.0e-03) NK-500 4.095e-02 (1.6e-03) $\dagger$ 4.280e-02 (1.6e-03) TSP-100 1.951e+03 (4.3e+01) $\dagger$ 2.090e+03 (3.1e+01) TSP-200 5.430e+03 (2.2e+02) $\dagger$ 6.044e+03 (2.3e+02) TSP-500 2.247e+04 (1.3e+03) $\dagger$ 2.620e+04 (1.2e+03) QAP-100 1.942e+15 (1.4e+14) 2.006e+15 (7.6e+13) QAP-200 1.499e+16 (6.2e+14) 1.506e+16 (7.9e+14) QAP-500 1.161e+17 (5.6e+15) $\dagger$ 1.255e+17 (6.7e+15)
@@
Table 4: The Hypervolume (Zitzler and Thiele 1999) results (mean and standard deviation) of SMS-EMOA-A (archive only to store non-dominated solutions) and SMS-EMOA-AR (archive with reuse) on KP, NK, TSP and QAP. The number next to each problem denotes the number of variables. The better mean is highlighted in bold.
@@ Appendix A Proof of Theorem 8 and Theorem 10
Definition 5 (OneMinMax (Giel and Lehre 2010)).
@@
Definition 6 (LeadingOnesTrailingZeroes (Laumanns, Thiele, and Zitzler 2004b)).
@@
Note that the analysis of Theorems 8 and 10 is very similar to that in (Bian et al. 2024). Although so, we still provide the proof in detail for completeness. The SMS-EMOA algorithm with archive reuse and one-point crossover operator is presented in Algorithm 2. The archive reuse mechanism is adapted such that each parent solution for recombination is selected with probability $1/2$ from the population and with probability $1/2$ from the archive. Note that one-point crossover selects a random point $i\in\{1,2,\ldots,n\}$ and exchanges the first $i$ bits of two parent solutions, which actually produces two new solutions, but the algorithm only picks the one that consists of the first part of the first parent solution and the second part of the second parent solution.

### v2

@@ Introduction
In this section, we conduct experiments on two artificial problems OneJumpZeroJump and OneJumpZeroJumpSS analyzed in this paper, and four well-known practical optimization problems: the multi-objective 0/1 knapsack (KP) (teghem_multi-objective_1994), travelling salesman problem (TSP) (ribeiro2002study), quadratic assignment problem (QAP) (knowles2003instance) and NK-landscapes (NK) (Aguirre2004), following the settings in (li_empirical_2024). Table 2 presents the results of SMS-EMOA solving OneJumpZeroJumpSS with problem size $n\in\{15,20,25,30\}$, $k=3$, and $a=2$, and Table 3 presents the results of SMS-EMOA soving OneJumpZeroJump with problem size $n\in\{15,20,25,30\}$ and $k=2$. We present the average number of generations over $1000$ runs for three variants: 1) SMS-EMOA with a large population size $\mu=n-2k+5$, 2) SMS-EMOA with an archive storing all non-dominated solutions ($\mu=5$), and 3) SMS-EMOA with archive reuse ($\mu=5$). For SMS-EMOA-A, since its expected running time on OneJumpZeroJump is at least $n^{\Omega(n)}$, we set the maximum number of iterations to $1,000,000$. For problem sizes $n=\{20,25,30\}$, a significant proportion of runs exceed this limit, thus these cases are marked with “-”. Results show that reusing the archive significantly reduces running time compared to using an archive only to store non-dominated solutions, and using a large population size.
@@
As for the four practical problems, each problem is instantiated at three sizes: 100, 200, and 500 variables. We conduct experiments using three SMS-EMOA variants: 1) With large population size ($\mu=1000$), 2) Archive without reuse ($\mu=200$), and 3) Archive with reuse ($\mu=200$). These variants follow common parameter settings: 1) the parent selection uses uniform random selection; 2) the crossover and mutation operators are problem-specific. For pseudo-Boolean problems (KP, NK), we use the uniform crossover (eiben2015introduction) with a rate of 1.0 and the bit-wise mutation (eiben2015introduction) with a rate of $1/n$, where $n$ is the number of variables. For TSP, we use the order-based crossover and the 2-opt mutation (eiben2015introduction). For QAP, we use the cycle-based crossover and the 2-swap mutation (eiben2015introduction). For both TSP and QAP, the crossover rate is 1.0 and the mutation rate is 0.05, following the practice in (li_empirical_2024). The budget for the algorithms is 10,000,000 fitness evaluations and the number of independent runs is set to 30. Table 4 presents the result of a widely used indicator, Hypervolume (HV) (zitzler1999multiobjective) of the three SMS-EMOA variants on the four problems. To compute the HV reference point, we sample 100,000 random solutions from the decision space and use the worst objective values among the non-dominated ones. Table 4 shows that reusing the archive can lead to a better mean HV across all problem instances.
@@
Figure 2 presents the non-dominated solutions obtained by the SMS-EMOA without archive reuse (black triangles), SMS-EMOA with archive reuse (blue diamonds), and SMS-EMOA with a large population size (yellow squares) on the multi-objective 0/1 Knapsack (KP), NK-Landscape (NK), TSP and QAP problems, with 100 (top panel), 200 (middle panel), and 500 (bottom panel) decision variables respectively. As can be seen, SMS-EMOA-RA (reusing the archive) consistently outperforms or matches SMS-EMOA (not reusing the archive) across all problem instances. The advantage of archive reuse becomes more pronounced as the problem size increases, particularly on the TSP.
[…]
@@
Table 4: Hypervolume (zitzler1999multiobjective) results (mean and standard deviation) of SMS-EMOAL (a large population size), SMS-EMOAA (archive without reuse), and SMS-EMOAAR (archive with reuse) on practical problems. The number in each problem name denotes the number of variables. The best mean per row is highlighted in bold.
@@ Appendix A Proof of Theorem 8 and Theorem 10
Definition 5 (OneMinMax (Giel10)).
@@
Definition 6 (LeadingOnesTrailingZeroes (LaumannsTEC04)).
@@
Note that the analysis of Theorems 8 and 10 is very similar to that in (bian2024archive). Although so, we still provide the proof in detail for completeness. The SMS-EMOA algorithm with archive reuse and one-point crossover operator is presented in Algorithm 2. The archive reuse mechanism is adapted such that each parent solution for recombination is selected with probability $1/2$ from the population and with probability $1/2$ from the archive. Note that one-point crossover selects a random point $i\in\{1,2,\ldots,n\}$ and exchanges the first $i$ bits of two parent solutions, which actually produces two new solutions, but the algorithm only picks the one that consists of the first part of the first parent solution and the second part of the second parent solution.